# Generative Diffusion Model Driven Massive Random Access in Massive MIMO Systems

### v1

@@ Abstract
Index Terms:
Deep learning, massive MIMO, massive random access, Transformer, generative models.
@@ I Introduction
To overcome the limitations of existing scheduling-based random access protocols, the grant-free non-orthogonal multiple access scheme has been proposed [1]. By eliminating the need for complex handshaking between users and the base station (BS), terminals can directly transmit uplink signals composed of non-orthogonal preamble sequences and payload data. This approach effectively reduces communication delays while enabling more devices to access the network simultaneously. However, it also increases the complexity of designing receiver detection algorithms. Researchers have developed various advanced receiver algorithms, which can be applied to one or more tasks in AUD, CE, and DD, depending on the specific challenges considered.
@@ I-A Related Work
The aforementioned research assumes that the channel between users and the BS is characterized by a multi-antenna Rayleigh channel, and it does not tackle the scenarios of non-Rayleigh channels. In non-Rayleigh environments, a common strategy involves the joint execution of AUD and CE tasks. This is often framed as a compressed sensing (CS) problem, which is solved using traditional greedy algorithms, Bayesian approaches, or deep learning methods. For example, the existing algorithms, such as orthogonal matching pursuit (OMP) [8], approximate message passing (AMP)[9], and orthogonal AMP (OAMP) [10], exploit the inherent sparsity of user activity to recover the user channel matrix while simultaneously identifying active users. To reduce computational complexity and enhance overall performance, model-driven deep learning approaches [11, 12] have been developed based on traditional CS algorithm frameworks. By integrating learnable parameters with the conventional backbones, such networks demonstrate superior performance under low pilot overhead conditions while significantly accelerating convergence. However, since the primary aim of CS is CE, with activity detection as a secondary outcome, joint AUD and CE methods typically necessitate higher pilot overhead than covariance-based techniques, which directly target the active user set detection [3].
Once the AUD and CE processes are completed, the massive MIMO DD problem at BS often becomes overdetermined. Various advanced detection algorithms have been developed to address this challenge. In addition to traditional methods like least squares (LS) and linear minimum mean square error (LMMSE), algorithms such as OAMP [13], expectation propagation (EP) [14], and deep learning-based techniques [15, 16, 17, 18] have demonstrated excellent performance. However, these DD methods rely on the channel state information (CSI) estimated in the earlier stage, and errors in CSI estimation will degrade the detection performance.
Performing joint CE and DD (JCEDD) can enhance overall achievable performance. The work [16] presents an iterative JCEDD scheme that accounts for channel estimation errors during data detection and employs a model-driven OAMP unfolding algorithm to improve detection. In the CE phase, demodulated data are used as pilots to augment LMMSE CE, resulting in superior performance. Additionally, message-passing algorithms have been developed for the joint AUD, CE, and DD [19, 20]. Despite their good performance, these algorithms necessitate intricate message-passing processes, with convergence and performance heavily reliant on initial estimates and the alignment between the assumed prior distribution and the underlying system’s probability distribution, thereby limiting their applicability.
@@ I-B Our Contributions
This paper proposes a new two-stage algorithm to solve the AUD, CE, and DD tasks in massive random access systems. The first stage identifies the active users set (AUS) through a Transformer-based detector, while the second stage utilizes generative diffusion models (GDMs) to perform JCEDD for active user equipment (UEs). By contrast, traditional approaches based on CS first explicitly estimate the channels and then perform AUD and DD, which degrades dramatically in the case of insufficient pilots, while covariance-based approaches only consider AUD and neglect CE and DD. Our main contributions are summarized as follows.
@@
  • To accommodate the distinct distribution characteristics of unknown continuous channel distributions and known discrete data constellation distributions, we propose an asynchronous iterative CE and DD framework. This framework integrates an LMMSE-based CE initialization method, which reduces the number of sampling steps and enhances convergence performance.
@@
The remainder of this paper is organized as follows. Section II briefly introduces the massive access system model, and Section II formulates the problems of AUD and JCEDD. Section IV details the proposed Transformer-based AUD scheme, while Section V elaborates on our generative models-based JCEDD scheme. Section VI presents the simulation results, and Section VII concludes the paper.
[…]
@@ III-A Stage 1-AUD
We assume that the original pilot sequence for each UE, denoted as $\tilde{\bm{p}}_{k}$, is selected from a large pilot pool $\tilde{\bm{P}}\!=\!\left[\tilde{\bm{p}}_{1},\tilde{\bm{p}}_{2},\dots,\tilde{\bm{p}}_{K}\right]\!\in\!\mathbb{C}^{L_{p}^{\text{max}}\times K}$ and is pre-allocated by the BS when the UE initially registers with the BS. The UE identity is then bound to this pilot sequence until it un-registers from the network. During the transmission procedure as shown in the left part of Fig. 1, only the first $L_{p}$ symbols of each pilot sequence are transmitted. That is, each pilot sequence is masked by a binary vector $\bm{m}\!=\!\left[1,\dots,1,0,\dots,0\right]^{\rm T}\!\in\!\left\{0,1\right\}^{L_{p}^{\text{max}}}$, whose first $L_{p}$ elements are ones and the remaining $\left(L_{p}^{\text{max}}-L_{p}\right)$ elements are zeros. Here, $L_{p}$ can take a value in the range $[L_{p}^{\text{min}},\,L_{p}^{\text{max}}]$. Let $\bm{M}\!=\!\bm{1}_{K}^{\rm T}\!\otimes\!\bm{m}\!\in\!\mathbb{B}^{L_{p}^{\text{max}}\times K}$. Then, at the AUD stage, the BS identifies the active UEs from the received pilot signals, expressed as:
@@
where $f_{\bm{\Phi}}\left(\tilde{\bm{Y}}_{p},\tilde{\bm{P}},L_{p}\right)$ is a neural network which maps the received signal and pilot sequences to $\hat{\bm{\alpha}}$, $\bm{\Phi}$ represents the parameters of the neural network, and $\mathcal{L}\left(\bm{\alpha},\hat{\bm{\alpha}}\right)$ is the loss function that quantifies the disparity between $\bm{\alpha}$ and $\hat{\bm{\alpha}}$.
@@ IV-A Preprocessing
Before inputting the user pilots and received signals into the Transformer-network block, preprocessing is required to ensure dimensional consistency. The user pilots need to be converted into real-valued representations, resulting in $K$ sequences, denoted as $\bm{z}_{k}^{o}\!\triangleq\!\big{[}\Re\{\tilde{\bm{p}}_{k}\}^{\rm T},\Im\{\tilde{\bm{p}}_{k}\}^{\rm T}\big{]}^{\rm T}\!\in\!\mathbb{R}^{2L_{p}^{\text{max}}}$, $1\!\leq\!k\!\leq\!K$, each of length $2L_{p}^{\text{max}}$. These sequences are then processed through a linear layer to produce the pilot feature sequences $\{\bm{z}_{k}^{1}\}_{k=1}^{K}$, each of length $d$. Concurrently, to handle varying pilot transmission lengths $L_{p}$ at the receiver, zero-padding is applied to the received signal $\bm{Y}_{p}\!\in\!\mathbb{C}^{L_{p}\times M}$ to convert it into $\tilde{\bm{Y}}_{p}\!\in\!\mathbb{C}^{L_{p}^{\text{max}}\times M}$ (see (3)). Subsequently, to leverage the multi-antenna observations at the BS, the covariance matrix [6, 4] of the received signals is calculated as $\bm{C}\!\triangleq\!\tilde{\bm{Y}}_{p}\tilde{\bm{Y}}_{p}^{\rm H}/M\!\in\!\mathbb{C}^{L_{p}^{\text{max}}\times L_{p}^{\text{max}}}$. This covariance matrix is transformed into a real-valued vectorized form, denoted as $\bm{z}_{K+1}^{o}\!\triangleq\!\big{[}\Re\{\text{vec}(\bm{C})\}^{\rm T},\Im\{\text{vec}(\bm{C})\}^{\rm T}\big{]}^{\rm T}\!\in\!\mathbb{R}^{2\left(L_{p}^{\text{max}}\right)^{2}}$. The resulting vector undergoes a linear layer transformation to achieve a dimensionality of $d$, denoted as $\bm{z}_{K+1}^{1}$. Finally, these $K$ pilot sequences and the received signal sequence, $\{\bm{z}_{k}^{1}\}_{k=1}^{K+1}$, are collectively fed into the subsequent heterogeneous Transformer for further feature extraction.
@@ IV-B Heterogeneous Transformer and Activity Decoder
Unlike the traditional Transformer, the heterogeneous Transformer introduced in [6] employs distinct sets of weight matrices for input embedding, multi-head attention (MHA), and the multi-layer perceptron to account for the different physical meanings of pilot sequences and received signals. Since the relative positional order of different sequences is not meaningful in our AUD task, we do not apply positional encoding as in conventional Transformers.
Assuming that the heterogeneous Transformer network consists of $N$ encoder layers, we conduct activity estimation for $K$ users based on the $K\!+\!1$ sequences output by the $N$-th encoder layer. This task is performed by the activity decoder, which computes the heterogeneous MHA between the final output feature sequence of the received signal $\bm{z}_{K+1}^{N+1}$ and pilot sequences $\{\bm{z}_{k}^{N+1}\}_{k=1}^{K}$. The resulting MHA output, $\bm{z}_{c}$, represents the synthesized contributions of all pilot sequences to the received signal. Next, the user activity is estimated as the weighted correlation between $\bm{z}_{c}$ and each pilot feature sequence given by:
@@
Here, $\{\lambda_{k}\}_{k=1}^{K}$ is the output of the activity decoder, $\bm{W}^{o}\!\in\!\mathbb{C}^{d\times d}$ is a learnable parameter matrix, and $C$ is a hyperparamter empirically set to $C\!=\!10$. User activity is then determined by a binary threshold detector: $\hat{\alpha}_{k}\!=\!\mathbb{I}(\lambda_{k}>0.5),\forall k$, where the indicator $\mathbb{I}(x)\!=\!1$ if $x$ is true, and 0 otherwise.
@@ IV-C Pilot Length Adaptive Module
Inspired by the attention feature module in [34], we propose a PLAM. By inserting a PLAM module between each pair of adjacent Transformer encoder layers, we provide additional pilot length information to the receiver. The structure of the proposed PLAM module is illustrated in Fig. 3. Its input consists of $(K\!+\!1)$ sequences from the output of the previous Transformer encoder layer, and its output comprises $(K\!+\!1)$ sequences of the same dimension, which serve as the input to the next layer. First, an average pooling operation is applied to the $(K\!+\!1)$ sequences along the sequence length dimension, producing a sequence of length $d$. This $d$-dimensional sequence is then concatenated with the pilot length information $L_{p}$ and processed through a two-layer fully connected network, generating a $d$-dimensional weight sequence. Finally, this weight sequence is used for position-wise dot multiplication with all input sequences, redistributing the weights of the input feature sequence and allowing the receiver to recognize the varying importance of different elements in the inputs.
@@ V-A Preliminary: Diffusion Process and SDE
Here, $p_{t}(\bm{x})$ denotes the probability of $\bm{x}$ at time $t$, and $\bar{\bm{w}}$ denotes the Wiener process when time flows backward from $T$ to 00. Existing generative models like SMLD [22] and DDPM [29] can be considered specific numerical discretization solvers of the SDE. For instance, in SMLD, the coefficients are $\bm{f}(\bm{x},t)\!=\!\bm{0}$ and $g(t)\!=\!\sqrt{\frac{\mathrm{d}(\sigma^{2}(t))}{\mathrm{d}t}}$, with $\sigma(t)$ serving as a predefined noise scheduling function. For DDPM, the drift and diffusion coefficients are given by $\bm{f}(\bm{x},t)\!=\!-\frac{\sigma(t)\bm{x}}{2}$ and $g(t)\!=\!\sqrt{\sigma(t)}$, respectively. By applying the discrete form of both the forward and reverse SDEs, we can derive the corresponding iteration equations used in SMLD and DDPM.

### v2

@@ I Introduction
To overcome the limitations of existing scheduling-based random access protocols, the grant-free non-orthogonal multiple access scheme has been proposed [1]. By eliminating the need for complex handshaking between users and the base station (BS), terminals can directly transmit uplink signals composed of non-orthogonal preamble sequences and payload data. This approach effectively reduces communication delays while enabling more devices to access the network simultaneously. However, it also increases the complexity of designing receiver detection algorithms. Previous studies have proposed several advanced receiver schemes, notably sourced random access and unsourced random access [3]. Sourced random access [4] employs pre-assigned preamble sequences to distinguish user identities and facilitates user activity detection and channel estimation using these preambles. In contrast, unsourced random access employs a common codebook, utilizing codewords to convey information and select coding, interleaving, and spreading techniques [5]. Moreover, reference [6] investigates a massive random access scheme utilizing individual user codebooks. This scheme recognizes the uniqueness of user codewords and encodes bits during the codeword selection process, integrating features from both sourced and unsourced random access mechanisms previously described. The reference [6] provides an in-depth performance analysis of massive random access under multiple-input multiple-output (MIMO) Rayleigh channels, taking into account the presence or absence of channel state information (CSI) at the receiver and whether the number of active users is known, thereby offering theoretical insights into performance limits. This paper mainly focuses on sourced random access schemes as in [4], tackling algorithm design challenges related to AUD, CE, and DD. For simplicity, we continue to refer to the discussed approach as massive random access.
@@ I-A Related Work
The aforementioned research assumes that the channel between users and the BS is characterized by a multi-antenna Rayleigh channel, and it does not tackle the scenarios of non-Rayleigh channels. In non-Rayleigh environments, a common strategy involves the joint execution of AUD and CE tasks. This is often framed as a compressed sensing (CS) problem, which is solved using traditional greedy algorithms, Bayesian approaches, or deep learning methods. For example, the existing algorithms, such as orthogonal matching pursuit (OMP) [12], approximate message passing (AMP)[13, 14], and orthogonal AMP (OAMP) [15], exploit the inherent sparsity of user activity to recover the user channel matrix while simultaneously identifying active users. To reduce computational complexity and enhance overall performance, model-driven deep learning approaches [16, 17] have been developed based on traditional CS algorithm frameworks. By integrating learnable parameters with the conventional backbones, such networks demonstrate superior performance under low pilot overhead conditions while significantly accelerating convergence. However, since the primary aim of CS is CE, with activity detection as a secondary outcome, joint AUD and CE methods typically necessitate higher pilot overhead than covariance-based techniques, which directly target the active user set detection [7].
Once the AUD and CE processes are completed, the massive MIMO DD problem at BS often becomes overdetermined. Various advanced detection algorithms have been developed to address this challenge. In addition to traditional methods like least squares (LS) and linear minimum mean square error (LMMSE), algorithms such as OAMP [18], expectation propagation (EP) [19], and deep learning-based techniques [20, 21, 22, 23] have demonstrated excellent performance. However, these DD methods rely on the CSI estimated in the earlier stage, and errors in CSI estimation will degrade the detection performance.
Performing joint CE and DD (JCEDD) can enhance overall achievable performance. The work [21] presents an iterative JCEDD scheme that accounts for channel estimation errors during data detection and employs a model-driven OAMP unfolding algorithm to improve detection. In the CE phase, demodulated data are used as pilots to augment LMMSE CE, resulting in superior performance. Additionally, several studies have addressed the joint AUD, CE, and DD problems, where message-passing algorithms or optimization-based methods have been developed to tackle these tasks [24, 25, 26]. Despite their strong performance, these algorithms require elaborate system configurations and dedicated algorithm design, which may not generalize effectively to practical systems or complex channel environments that are challenging to model, thus limiting their applicability. Consequently, they are not included in the comparisons of this paper.
@@ I-B Our Contributions
This paper proposes a new two-stage algorithm to solve the AUD, CE, and DD tasks in massive random access systems. The first stage identifies the active users set (AUS) through a Transformer-based detector, while the second stage utilizes generative diffusion models (GDMs) to perform JCEDD for active user equipment (UEs). The existing Transformer-based deep learning AUD scheme [10] fails to effectively address the challenges posed by variable pilot lengths in massive random access problems. Furthermore, the performance of current JCEDD schemes relying on ALD-based generative methods [33] has not been comprehensively evaluated and is susceptible to converging to local optima. To tackle these issues, our research makes the following key contributions:
@@
  • To accommodate the distinct characteristics of unknown continuous channel distributions and known discrete data constellation distributions, we propose an asynchronous iterative CE and DD framework. This framework integrates an LMMSE-based CE initialization method, which reduces the number of sampling steps and enhances convergence performance.
@@
The remainder of this paper is organized as follows. Section II briefly introduces the massive access system model, and Section III formulates the problems of AUD and JCEDD. Section IV details the proposed Transformer-based AUD scheme, while Section V elaborates on our generative models-based JCEDD scheme. Section VI presents the simulation results, and Section VII concludes the paper.
Notations: Matrices and column vectors are denoted by uppercase and lowercase boldface letters, respectively. $\bm{I}_{N}$ is the $N\times N$ identity matrix, $\bm{0}_{N\times M}$ is the $N\times M$ all-zero matrix, and $\bm{1}_{N}$ is the $N\times 1$ all-one vector. $(\cdot)^{\rm T}$, $(\cdot)^{*}$, $(\cdot)^{\rm H}$, $(\cdot)^{-1}$, and $\mathbb{E}\{\cdot\}$ denote the transpose, conjugate, Hermitian transpose, inversion, and expectation operations, respectively. $|\mathcal{A}|_{c}$ indicates the cardinality of the set $\mathcal{A}$. $\mathrm{diag}(\bm{a})$ transforms vector $\bm{a}$ into the corresponding diagonal matrix. $\mathcal{N}\left(\bm{x};\bm{a},\bm{A}\right)$ denotes the probability density function (PDF) of a Gaussian random vector $\bm{x}$ with mean $\bm{a}$ and covariance matrix $\bm{A}$, and the PDF of a complex-valued Gaussian random vector is denoted by $\mathcal{N}_{c}(\cdot;\cdot,\cdot)$. $\bm{X}_{[:,\mathcal{A}]}$ ($\bm{X}_{[\mathcal{A},:]}$) refers to the sub-matrix comprising the columns (rows) of $\bm{X}$ identified by the index set $\mathcal{A}$, while $\otimes$ and $\odot$ denotes Kronecker and Hadamard product operations, respectively. The operator $\text{vec}(\cdot)$ stacks a matrix column by column into a vector, and $\bm{X}]_{i,j}$ is the $(i,j)$-th element of $\bm{X}$. $\left\lfloor\cdot\right\rfloor$ denotes the integer floor operator.
@@ III-A Stage 1-AUD
We assume that the original pilot sequence for each UE, denoted as $\tilde{\bm{p}}_{k}$, is selected from a large pilot pool $\tilde{\bm{P}}\!=\!\left[\tilde{\bm{p}}_{1},\tilde{\bm{p}}_{2},\dots,\tilde{\bm{p}}_{K}\right]\!\in\!\mathbb{C}^{L_{p}^{\text{max}}\times K}$ and is pre-allocated by the BS when the UE initially registers with the BS. The UE identity is then bound to this pilot sequence until it un-registers from the network. During the transmission procedure as shown in the left part of Fig. 1, only the first $L_{p}$ symbols of each pilot sequence are transmitted. That is, each pilot sequence is masked by a binary vector $\bm{m}\!=\!\left[1,\dots,1,0,\dots,0\right]^{\rm T}\!\in\!\left\{0,1\right\}^{L_{p}^{\text{max}}}$, whose first $L_{p}$ elements are ones and the remaining $\left(L_{p}^{\text{max}}-L_{p}\right)$ elements are zeros. Here, $L_{p}$ can take a value in the range $L_{p}^{\text{min}},\,L_{p}^{\text{max}}]$. Let $\bm{M}\!=\!\bm{1}_{K}^{\rm T}\!\otimes\!\bm{m}\!\in\!\mathbb{B}^{L_{p}^{\text{max}}\times K}$. Then, at the AUD stage, the BS identifies the active UEs from the received pilot signals, expressed as:
@@
[…]
@@ IV-A Preprocessing
Before inputting the user pilots and received signals into the Transformer-network block, preprocessing is required to ensure dimensional consistency. The user pilots need to be converted into real-valued representations, resulting in $K$ sequences, denoted as $\tilde{\bm{z}}_{k}\!\triangleq\!\big[\Re\{\tilde{\bm{p}}_{k}\}^{\rm T},\Im\{\tilde{\bm{p}}_{k}\}^{\rm T}\big]^{\rm T}\!\in\!\mathbb{R}^{2L_{p}^{\text{max}}}$, $1\!\leq\!k\!\leq\!K$, each of length $2L_{p}^{\text{max}}$. These sequences are then processed through a linear layer to produce the pilot feature sequences $\{\bm{z}_{k}^{0}\}_{k=1}^{K}$, each of length $d$. Concurrently, to handle varying pilot transmission lengths $L_{p}$ at the receiver, zero-padding is applied to the received signal $\bm{Y}_{p}\!\in\!\mathbb{C}^{L_{p}\times M}$ to convert it into $\tilde{\bm{Y}}_{p}\!\in\!\mathbb{C}^{L_{p}^{\text{max}}\times M}$ (see (3)). Subsequently, to leverage the multi-antenna observations at the BS, the covariance matrix [10, 8] of the received signals is calculated as $\bm{C}\!\triangleq\!\tilde{\bm{Y}}_{p}\tilde{\bm{Y}}_{p}^{\rm H}/M\!\in\!\mathbb{C}^{L_{p}^{\text{max}}\times L_{p}^{\text{max}}}$. This covariance matrix is transformed into a real-valued vectorized form, denoted as $\tilde{\bm{z}}_{K+1}\!\triangleq\!\big[\Re\{\text{vec}(\bm{C})\}^{\rm T},\Im\{\text{vec}(\bm{C})\}^{\rm T}\big]^{\rm T}\!\in\!\mathbb{R}^{2\left(L_{p}^{\text{max}}\right)^{2}}$. The resulting vector undergoes a linear layer transformation to achieve a dimensionality of $d$, denoted as $\bm{z}_{K+1}^{0}$. Finally, these $K$ pilot sequences and the received signal sequence, $\{\bm{z}_{k}^{0}\}_{k=1}^{K+1}$, are collectively fed into the subsequent heterogeneous Transformer for further feature extraction.
@@ IV-B Heterogeneous Transformer and Activity Decoder
Unlike the traditional Transformer, the heterogeneous Transformer introduced in [10] employs distinct sets of weight matrices for input embedding, multi-head attention (MHA), and the feed-forward network (FFN) to account for the different physical meanings of pilot sequences and received signals. To facilitate understanding, we illustrate the heterogeneous Transformer structure and MHA in Appendix A. Since the relative positional order of different sequences is not meaningful in our AUD task, we do not apply positional encoding as in conventional Transformers.
Assuming that the heterogeneous Transformer network consists of $N$ encoder layers, we conduct activity estimation for $K$ users based on the K+1K\!+\!1 sequences output by the $N$-th encoder layer. This task is performed by the activity decoder, which computes the heterogeneous MHA between the final output feature sequence of the received signal $\bm{z}_{K+1}^{N+1}$ and pilot sequences $\{\bm{z}_{k}^{N+1}\}_{k=1}^{K}$. The resulting MHA output, $\bm{z}_{c}$, represents the synthesized contributions of all pilot sequences to the received signal. Next, the user activity is estimated as the weighted correlation between $\bm{z}_{c}$ and each pilot feature sequence. It is given by:
@@
Here, $\{\lambda_{k}\}_{k=1}^{K}$ is the output of the activity decoder, $\bm{W}^{o}\!\in\!\mathbb{C}^{d\times d}$ is a learnable parameter matrix, and $C$ is a hyperparamter empirically set to C=10C\!=\!10. User activity is then determined by a binary threshold detector: $\hat{\alpha}_{k}\!=\!\mathbb{I}(\lambda_{k}>0.5),\forall k$, where the indicator $\mathbb{I}(x)\!=\!1$ if $x$ is true, and 0 otherwise.
@@ IV-C Pilot Length Adaptive Module
Inspired by the attention feature module in [40], we propose a PLAM. By inserting a PLAM module between each pair of adjacent Transformer encoder layers, we provide additional pilot length information to the receiver. The structure of the proposed PLAM module is illustrated in Fig. 3. Its input consists of (K+1)(K\!+\!1) sequences from the output of the previous Transformer encoder layer, and its output comprises (K+1)(K\!+\!1) sequences of the same dimension, which serve as the input to the next layer. First, an average pooling operation is applied to the (K+1)(K\!+\!1) sequences along the sequence length dimension, producing a sequence of length $d$. This $d$-dimensional sequence is then concatenated with the pilot length information $L_{p}$ and processed through a two-layer fully connected network, generating a $d$-dimensional weight sequence. Finally, this weight sequence is used for position-wise dot multiplication with all input sequences, redistributing the weights of the input feature sequence and allowing the receiver to recognize the varying importance of different elements in the inputs.
@@ V-A Preliminary: Diffusion Process and SDE
Here, $p_{t}(\bm{x})$ denotes the probability of $\bm{x}$ at time $t$, and $\bar{\bm{w}}$ denotes the Wiener process when time flows backward from $T$ to 0. Existing generative models like SMLD [28] and DDPM [35] can be considered specific numerical discretization solvers of the SDE. For instance, in SMLD, the coefficients are $\bm{f}(\bm{x},t)\!=\!\bm{0}$ and $g(t)\!=\!\sqrt{\frac{\mathrm{d}(\sigma^{2}(t))}{\mathrm{d}t}}$, with $\sigma(t)$ serving as a predefined noise scheduling function. For DDPM, the drift and diffusion coefficients are given by $\bm{f}(\bm{x},t)\!=\!-\frac{\sigma(t)\bm{x}}{2}$ and $g(t)\!=\!\sqrt{\sigma(t)}$, respectively. By applying the discrete form of both the forward and reverse SDEs, we can derive the corresponding iteration equations used in SMLD and DDPM.